Part 1. If the absolute pressure in a tank is 140 kPa, determine the pressure head in mm of mercury. the atmospheric pressure is 100 kPa.


Part 2. Oxygen in a tank has an absolute pressure of 130 kPa. Determine the pressure head in mm of mercury. The atmospheric pressure is 102 kPa.



1
Expert's answer
2021-11-21T17:27:35-0500

1)

"p'=p-p_0=40~\text{kPa}=300" mm of mercury.

2)

"p'=p-p_0=28~\text{kPa}=210" mm of mercury.